OpenStudy (anonymous):

Equation: y=3x+5 a) given that dx/dt = 2, find dy/dy when x=1. b)given that dy/dt = -1, find dx/dt when x=0.

OpenStudy (anonymous):

a) differntiate with respect to t dy/dt = 3 (dx/dt) dx/dt = 2 so dy/dt is 6. b) dy/dt =3(dx/dt) so dx/dt = 1/3 * (dy/dt) now dy/dt = -1 so dx/dt is -1/3